Aqual:

Let's say you flip two coins, roll a six-sided die, and pick a random card out of a standard deck of cards. What is the probability of landing on two tails, rolling on an odd number, and drawing a club card?

Aqual:

@Ultrilliam

MiraAngel:

thx

Aqual:

@silvernight269

MiraAngel:

@MrMudd183

Aqual:

@Hero

Salmon:

@Shadow

Shadow:

Probability of two tails: \[\frac{ 1 }{2 } \times \frac{ 1 }{ 2 } = ?\] Probability of an odd number from a six sided die: (1,3,5), there are three odd numbers on a six sided die \[\frac{ 3 }{ 6 } \rightarrow ?\] Simplify Probability of a club card from a deck: Googled and there are apparently 13 cards in a deck, which is 52 total cards \[\frac{ 13 }{ 52 }\]

Shadow:

Multiply the probabilities together and you should have the probability of all the events occurring.

Shadow:

Make sure to do the final steps of each part, such as finishing the multiplication and simplifying the fractions. Should be easy.

Shadow:

^ This assumes that they're asking for the probability of all these events occurring together. If they want just the probability of each of the three events, don't multiply the probabilities together.

MiraAngel:

Thank youuuuuuuu!